OpenStudy (anonymous):

Select the word that best completes the statement. The sum of two integers is _____ positive. (1 point) always sometimes or never (hold on give me a minute to think about my answer)

OpenStudy (anonymous):

sometimes

OpenStudy (anonymous):

sometimes... it's exactly half the time seeing as an integer can be positive or negative and extend to infinity in both directions lol. Example: -1+-1=-2 but 1+1=2

OpenStudy (anonymous):

On a side note is your avatar from Black Rock Shooter?

OpenStudy (anonymous):

yes @EllaRu

OpenStudy (anonymous):

Yay! Good anime. :)

OpenStudy (anonymous):

yes ^^ @EllaRu is the sum of this 7? 16 + ( –9 ) =

OpenStudy (anonymous):

yup. 16 + (-9) is the same as 16-9 which =7